OpenStudy (anonymous):

Please Help!? divide: x-2 divided into x^3 + 8

OpenStudy (aravindg):

are you sure its x^3+8 ?

OpenStudy (anonymous):

im sorry minus 8 typo (:

OpenStudy (anonymous):

your problem is = [x^3-8]/(x-2) =(x^3-2^3)/(x-2) =(x-2)(x^2+2x+2^2)/(x-2) =x^2+2x+4

OpenStudy (anonymous):

Or you can use synthetic division and come up to the answer

OpenStudy (anonymous):

can you show me the synthetic division way? please and thank you

OpenStudy (anonymous):

Well actually that would be too long but I can direct you to a site that can assist you by showing you examples

OpenStudy (anonymous):

Yes please!!! my instructor told the class to try that way but im so unsure of how to divide with the x^3 - 8 ?

OpenStudy (anonymous):

Try which way?

OpenStudy (anonymous):

the synthetic division way (: sorry

OpenStudy (anonymous):

http://www.purplemath.com/modules/synthdiv.htm Ok here is the site and it will walk you straight through I hope you do well :)

OpenStudy (anonymous):

Thank you so much!!!!!!!!!!!!! (: